OpenStudy (anonymous):

find the slope of a line parallel to x-y=16

OpenStudy (anonymous):

1

OpenStudy (lalaly):

re arrange -y=16-x y=x-16 slope =1

OpenStudy (anonymous):

y=mx +b ,m is the slope and parrelel lines have same slopes